# Scaling Offline Model-Based RL via Jointly-Optimized World-Action Model Pretraining

### v3

@@ 2 Related work
Offline RL algorithms learn a policy entirely from the static offline dataset without online interactions. Model-free offline RL incorporates conservatism to mitigate extrapolation error (Jin et al., 2021) primarily through policy constraints (Fujimoto et al., 2019; Kumar et al., 2019; Fujimoto & Gu, 2021; Cheng et al., 2024; Liu et al., 2024; 2025) and value regularization (Kumar et al., 2020b). Model-based offline RL approximates the environment using world models and performs conservative policy optimization (Yu et al., 2020; 2021). While these works focus on single-task settings, our work explores scaling offline model-based RL across diverse, challenging multi-task Atari games (Lee et al., 2022; Kumar et al., 2023; Wu et al., 2024) aiming for sample-efficient transfer to novel games.
@@ 3.1 Online distributional rl (C51)
[…]
@@ 4.3 Training pipeline
We employ this two-stage training approach to stabilize and accelerate the overall training process. In our pretraining experiments, we set $M_{1}=250$k and $M_{2}=1.5$M, totaling $1.75$M gradient steps.
@@ 5.3 How does JOWA perform on multi-games in low-data regime?
Assault 222.4 742.0 1203.5 1741.5 1915.5 2292 925.9 1578.3 1733.9 2302
Atlantis 12850 29028.1 37812.5 2565750 2108166.7 49100 86250 41662.5 570862.5 2690387.5
BeamRider 363.9 16926.5 786 6011.3 4149.8 7023.5 6822 880.3 2547.4 3498
@@
DemonAttack 152.1 1971 2611.3 6549.4 3455.7 7936.5 5806.4 1233.8 4584.4 3547.8
NameThisGame 2292.3 8049 6045 6610.5 7060 6461.7 6236 7335.6 12706.9 11421
Phoenix 761.4 7242.6 1446.9 5120.5 5320 4961.7 3814.5 1608.8 5065 5348
Seaquest 68.4 42054.7 120 2720 3160.4 650 1760 1033.1 1490 2725
SpaceInvaders 148 1668.7 605 742.5 513.8 714.4 641.2 694.7 969.1 744.7
StarGunner 664.0 10250 1493.8 8625.5 9550 4728.6 4936.4 5737.5 21231.3 18150
TimePilot 3568 5229.2 762.5 3866.7 2812.5 4072.7 4166.7 3268.8 3831.3 3669
Zaxxon 32.5 9173.3 0 462.5 325.5 0 312.5 0 225 2163
@@ 5.6 Ablation Study
To further demonstrate the efficiency of our planning algorithm, we compare it with MCTS. Because Muzero-style (Schrittwieser et al., 2020) MCTS requires access to the $V$-function and policy networks $\pi$ while JOWA only estimates the optimal $Q$-function, we use $\max_{a}Q(s,a)$ as the $V$-value $V(s)$ and use an energy-based policy to compute the action probability, i.e., $\pi(\cdot|s)=\operatorname{softmax}(Q(s,\cdot)/t)$, where $t$ is the temperature. We conduct a grid search on the implementation choices and hyperparameters for MCTS and show the details in Appendix F. We report the frames per second (FPS) and HNS on 7 games, where bare JOWA-150M performs poorly, in Table 5. Observe that ours not only is 10$\times$ faster than MCTS but also exceeds MCTS by 71.0% on Mean HNS. Moreover, we empirically observe that MCTS is highly sensitive to the temperature and max depth of the tree while the long execution time makes it inconvenient to tune hyperparameters.
@@
By default, we do not employ planning or synthetic data during pretraining due to the significant increase in training time (approximately 10$\times$ slower, extending training to months). For this ablation study, we enable synthetic data in pretraining for JOWA-40M and train on 6 games for 1M steps, which takes 31 days. Specifically, we sample batch of 4-step segments and interact in the imagined MDP with $\epsilon$-greedy policy to synthesis the last 4 steps. Then we optimize JOWA with half real data and half synthetic data using COMBO loss as $\mathcal{L}_{\text{action}}$. The results are shown in the last main column of Table 6.
@@ Appendix B Derivation from Equation (10) to Equation (11)
However, we can use the conclusion derived with ground-truth functions to make assumptions to reduce the policy space for search. We assume that the optimal policy for formula (17) maps states to actions with top-K highest Q-values. Denote the constrained policy space as $\Pi_{\hat{Q}^{*}}$, where $\forall\pi\in\Pi_{\hat{Q}^{*}},\forall s\in\mathcal{S},\forall a\notin\text{top-}K(\hat{Q}^{*}(s,\cdot))$, we have $\pi(a|s)=0$. we make the following assumption: the optimal policy of formula (17) is in the constrained policy space, i.e.,$\pi^{*}_{\hat{Q}^{*}}\in\Pi_{\hat{Q}^{*}}$. Under this assumption, formula (17) is equivalent to:
@@ Appendix F More experiments and Raw scores
Capability emergence refers to the phenomenon where large models suddenly exhibit new abilities or significantly enhanced performance, typically occurring after reaching certain scale thresholds in model size. For example, JOWA-150M achieves a significantly higher score than JOWA-70M and other baselines on Zaxxon, which seems like the “emergent” phenomenon. However, after examining the reward of Zaxxon, we conclude that this “emergent behavior” is actually attributable to the nonlinear reward function. In addition to the common rewards of around 100, Zaxxon also defines a huge reward of nearly 5000. JOWA-70M gets this huge reward in 1 of the 16 rollouts while JOWA-150M gets it in 7 rollouts during evaluation. When we clip the rewards to the range $[-1,1]$, JOWA-70M and JOWA-150M achieve scores of 0.08 and 0.44 respectively, indicating no true “emergent” phenomenon. Schaeffer et al. (2024) draws a similar conclusion that “nonlinear or discontinuous metrics cause the emergent abilities”.

### v4

@@ 2 Related work
Offline RL algorithms learn a policy entirely from the static offline dataset without online interactions. Model-free offline RL incorporates conservatism to mitigate extrapolation error (Jin et al., 2021) primarily through policy constraints (Fujimoto et al., 2019; Kumar et al., 2019; Fujimoto and Gu, 2021; Cheng et al., 2024; Liu et al., 2024; 2025) and value regularization (Kumar et al., 2020b). Model-based offline RL approximates the environment using world models and performs conservative policy optimization (Yu et al., 2020; 2021). While these works focus on single-task settings, our work explores scaling offline model-based RL across diverse, challenging multi-task Atari games (Lee et al., 2022; Kumar et al., 2023; Wu et al., 2024) aiming for sample-efficient transfer to novel games.
@@ 3.1 Online distributional rl (C51)
$\mathcal{T}^{*}Z(s,a):\overset{D}{=}R(s,a)+\gamma Z(s^{\prime},\arg\max_{a^{\prime}}Q(s^{\prime},a^{\prime}))$ (1)
where the formula $Y:\overset{D}{=}U$ denotes equality of probability laws, that is the random variable $Y$ is distributed according to the same law as $U$. The C51 algorithm (Bellemare et al., 2017) models $Z(s,a)$ using a discrete distribution supported on $N$ fixed atoms $z_{1}\leq\cdots\leq z_{N}$ uniformly spaced over a predetermined interval. Given a current value distribution, C51 applies a projection step to map the target distribution onto its finite element support and optimizes as follows:
@@ 4.3 Training pipeline
We employ this two-stage training approach to stabilize and accelerate the overall training process. In our pretraining experiments, we set $M_{1}=250$k and $M_{2}=1.5$M, totaling 1.751.75M gradient steps.
@@ 5.3 How does JOWA perform on multi-games in low-data regime?
Assault 222.4 742.0 1203.5 1741.5 1915.5 2292 925.9 1578.3 1402.8 2302
Atlantis 12850 29028.1 37812.5 2565750 2108166.7 49100 86250 41662.5 228825 2690387.5
BeamRider 363.9 16926.5 786 6011.3 4149.8 7023.5 6822 880.3 854.4 3498
@@
DemonAttack 152.1 1971 2611.3 6549.4 3455.7 7936.5 5806.4 1233.8 1409.4 3547.8
NameThisGame 2292.3 8049 6045 6610.5 7060 6461.7 6236 7335.6 7507.1 11421
Phoenix 761.4 7242.6 1446.9 5120.5 5320 4961.7 3814.5 1608.8 5070 5348
Seaquest 68.4 42054.7 120 2720 3160.4 650 1760 1033.1 1040 2725
SpaceInvaders 148 1668.7 605 742.5 513.8 714.4 641.2 694.7 805.3 744.7
StarGunner 664.0 10250 1493.8 8625.5 9550 4728.6 4936.4 5737.5 21700 18150
TimePilot 3568 5229.2 762.5 3866.7 2812.5 4072.7 4166.7 3268.8 1100.3 3669
Zaxxon 32.5 9173.3 0 462.5 325.5 0 312.5 0 200 2163
@@ 5.6 Ablation Study
[…]
@@
[…]
@@ Appendix B Derivation from Equation (10) to Equation (11)
[…]
@@ Appendix F More experiments and Raw scores
Capability emergence refers to the phenomenon where large models suddenly exhibit new abilities or significantly enhanced performance, typically occurring after reaching certain scale thresholds in model size. For example, JOWA-150M achieves a significantly higher score than JOWA-70M and other baselines on Zaxxon, which seems like the “emergent” phenomenon. However, after examining the reward of Zaxxon, we conclude that this “emergent behavior” is actually attributable to the nonlinear reward function. In addition to the common rewards of around 100, Zaxxon also defines a huge reward of nearly 5000. JOWA-70M gets this huge reward in 1 of the 16 rollouts while JOWA-150M gets it in 7 rollouts during evaluation. When we clip the rewards to the range [1,1][-1,1], JOWA-70M and JOWA-150M achieve scores of 0.08 and 0.44 respectively, indicating no true “emergent” phenomenon. Schaeffer et al. (2024) draws a similar conclusion that “nonlinear or discontinuous metrics cause the emergent abilities”.